OpenStudy (anonymous):

A' and ' B ' can do a piece of work in 25 days and 30 days respectively. Both start the work together but ' A' leaves the work 8 days before its completion. Find the time in which the work is finished.

OpenStudy (shubhamsrg):

whats your attempt(progress) ?

OpenStudy (anonymous):

1/25+1/30=a+b 's one days work (1-11/150)=(x-8)

OpenStudy (shubhamsrg):

A does 1/25 of the work in 1 day B does 1/30 of the work in 1 day together, in 1 day they do 1/25 + 1/30 of the work Let they both work together for x days then, they'll do x/25 + x/30 of the job after that, B does rest of the job alone for next 8 days => 8/30 of the work hence x/25 + x/30 = 1 -(8/30) = 22/30 solve for "x+8" . Hope it helps?

OpenStudy (anonymous):

10 days.